A man of 30 kg walking on a plank of 100kg find the displacement of plank?
1
Expert's answer
2020-09-21T08:28:46-0400

Let plank is large enough. Since external force acting on the plank is zero hence center of mass of the system will remain same.

If d is distance travelled by man and x is distance travelled by plank then in center of mass frame,

"30 d = (30 + 100)(x)"


"x = \frac{30}{130} d = \frac{3}{13}d"